The mean of a data set is µ = 42 and the closest outlier above the mean is 63. For this to be true, the standard deviation of the data set must be _______________.
Answer #1

Answer

we know that any value which is away from 3 standard deviation from the mean, then it is considered as an outlier.

we have mean = 42 and outlier = 63

Suppose standard deviation is x

then we can write it as

outlier = mean +(3*standard deviation)

setting the given values, we get

63 = 42 + 3x

subtracting 42 from each side, we get

63-42 = 3x

or 21 = 3x

dividing both sides by 3, we get

x = 21/3 = 7

So, standard deviation is 7